# HRSeg: High-Resolution Visual Perception and Enhancement for Reasoning Segmentation

### v1

@@ Abstract.
The reasoning segmentation task involves segmenting objects within an image by interpreting implicit user instructions, which may encompass subtleties such as contextual cues and open-world knowledge. Despite significant advancements made by existing approaches, they remain constrained by low perceptual resolution, as visual encoders are typically pre-trained at lower resolutions. Furthermore, simply interpolating the positional embeddings of visual encoders to enhance perceptual resolution yields only marginal performance improvements while incurring substantial computational costs. To address this, we propose HRSeg, an efficient model with high-resolution fine-grained perception. It features two key innovations: High-Resolution Perception (HRP) and High-Resolution Enhancement (HRE). The HRP module processes high-resolution images through cropping, integrating local and global features for multi-granularity quality. The HRE module enhances mask features by integrating fine-grained information from high-resolution images, refining their alignment with text features for precise segmentation. Extensive ablation studies validate the effectiveness of our modules, while comprehensive experiments on multiple benchmark datasets demonstrate HRSeg’s superior performance. Code will be available at https://github.com/WeihuangLin/HRSeg.
Reasoning Segmentation, Multimodal Large Language Models, High-resolution Perception
copyright: acmcopyrightjournalyear: 2025isbn: conference: Proceedings of the 33th ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandccs: Computing methodologies Image segmentation
@@ 3.3. High-Resolution Enhancement
Method val test
overall short query long query overall
gIoU cIoU gIoU cIoU gIoU cIoU gIoU cIoU
OVSeg (Liang et al., 2023) 28.5 18.6 18.0 15.5 28.7 22.5 26.1 20.8
GRES (Liu et al., 2023a) 22.4 19.9 17.6 15.0 22.6 23.8 21.3 22.0
X-Decoder (Zou et al., 2023) 22.6 17.9 20.4 11.6 22.2 17.5 21.7 16.3
SEEM (Zou et al., 2024) 25.5 21.2 20.1 11.5 25.6 20.8 24.3 18.7
LISA (Lai et al., 2024) 44.4 46.0 37.6 34.4 36.6 34.7 36.8 34.1
LLM-Seg (Wang and Ke, 2024) 47.4 35.4 39.8 32.9 41.3 33.8 40.2 33.4
SAM4MLLM (Chen et al., 2024a) 46.7 48.1 - - - - - -
HRSeg 50.1 47.1 42.1 37.7 47.9 43.7 46.6 42.6
HRSeg-LLaVA-v1.6 56.7 55.4 47.9 44.2 56.6 53.2 54.6 51.2
LISA (ft) 51.0 50.6 40.6 40.6 49.4 51.0 47.3 48.4
LLM-Seg (ft) 52.3 47.5 41.1 40.2 49.8 49.1 47.9 46.2
HRSeg (ft) 57.4 54.7 43.4 41.7 54.4 54.5 51.6 50.5
HRSeg-LLaVA-v1.6 (ft) 64.9 63.1 49.5 48.7 59.2 58.9 57.0 57.2
@@ 4.2. Implementation Details
We trained HRSeg for one day on 2 NVIDIA 3090 GPUs using the DeepSpeed engine. The mask proposals are generated using a $32\times 32$ grid of points to prompt SAM. The default MLLM is LLaVA-7B-v1-1, with LoRA applied for efficient fine-tuning. The image encoder used for the HRP module is DINOv2-ViT-L (Oquab et al., 2023), which is pretrained on images of size $518\times 518$ pixels. We use AdamW (Loshchilov, 2017) as the optimizer and apply WarmupDecayLR to adjust the learning rate. The initial learning rate is set to 0.0003. We use a batch size of 2 per GPU, with a gradient accumulation step of 10. For high-resolution images, we set $N=2$ by default, which corresponds to an image size of $518\times 2=1036$ The $\gamma$ in the HRP module is set to 0.8 to weight the features, and the $L$ value is set to 2 in the HRE module. For the loss function, the weights for the text loss ($\lambda_{text}$) and selection loss ($\lambda_{sel}$) are both set to 1.0.
@@ 4.3. Results on ReasonSeg
Table 1 provides a detailed comparison of our proposed HRSeg against various state-of-the-art segmentation models. Notably, HRSeg is a novel model tailored for high-resolution reasoning segmentation tasks.
The experimental results reveal that HRSeg consistently outperforms prior methods across all metrics and evaluation splits on the ReasonSeg dataset. Without fine-tuning, HRSeg achieves remarkable gains in gIoU and cIoU, outperforming LLM-Seg by 2.7% and 8.8% on the validation set, and by 6.4% and 9.2% on the overall test set, respectively. Furthermore, HRSeg demonstrates superior handling of both short and long queries, achieving a substantial 5.6% and 11.0% improvement in gIoU and cIoU over LLM-Seg for short queries, and a 6.6% and 9.9% improvement for long queries. Fine-tuning on the ReasonSeg training split further boosts the performance of HRSeg, achieving superior results across all evaluation metrics.
The experiments also reveal that in long-query instructions, which are more complex and require richer visual details, HRSeg outperforms LLM-Seg by 6.6%. This improvement is 4.3% higher than that observed in short-query scenarios. This demonstrates that HRSeg excels at capturing fine-grained visual features, allowing it to generate high-quality mask features and achieve better alignment with complex queries.
Therefore, we attribute our model’s superior performance to its ability to leverage high-resolution visual information, which not only enhances the level of detail in the segmentation but also facilitates the handling of implicit query text, thereby improving overall segmentation accuracy.
@@ 4.5. Results on Referring Segmentation
Method refCOCO refCOCO+ refCOCOg
val testA testB val testA testB val(U) test(U)
traditional methods
MCN (Luo et al., 2020) 62.4 64.2 59.7 50.6 55.0 44.7 49.2 49.4
VLT (Ding et al., 2021) 67.5 70.5 65.2 56.3 61.0 50.1 55.0 57.7
CRIS (Wang et al., 2022) 70.5 73.2 66.1 62.3 68.1 53.7 59.9 60.4
LAVT (Yang et al., 2022) 72.7 75.8 68.8 62.1 68.4 55.1 61.2 62.1
ReLA (Liu et al., 2023a) 73.8 76.5 70.2 66.0 71.0 57.7 65.0 66.0
SEEM (Zou et al., 2024) - - - - - - 65.7 -
LLM based
HRSeg-LLaVA-v1 77.1 79.7 73.2 66.4 72.4 58.9 71.4 72.8
HRSeg-LLaVA-v1.6 81.2 83.4 78.1 73.7 79.1 67.1 75.6 76.4
@@ 4.6. Ablation Study
Study on the weight $\gamma$. To evaluate the weight allocation when combining global features and region features, we analyze the impact of varying the value of $\gamma$ within the HRP module. As shown in Table 5, the results demonstrate that the choice of $\gamma$ significantly influences the model’s ability to capture hybrid information from these two types of features. When $\gamma$ is too low, the contribution of global features is diminished, limiting the model’s holistic understanding of the scene. Conversely, overly high values of $\gamma$ reduce the impact of region-specific details, which are crucial for precise reasoning segmentation. The optimal setting, $\gamma=0.8$, achieves the best balance, effectively leveraging both global context and detailed regional information to enhance the model’s reasoning capability.
@@
Study on the image resolution. To further investigate the impact of increased image resolution on the performance of our model, we conducted an ablation study on image resolution. As shown in Table 6, $N$ represents the magnification factor of the pre-trained encoder resolution. The results indicate that increasing the resolution of the input images enhances the model’s ability to capture finer details, which leads to better segmentation performance. Specifically, higher resolution allows the model to preserve more granular information, thus improving both global and local reasoning. However, this performance gain gradually diminishes as the resolution increases. While performance continues to improve with higher values of $N$, the computational cost associated with larger image sizes becomes significantly higher, leading to increased memory usage and processing time. Thus, $N=2$ was chosen as the optimal resolution for our model, as the marginal gains beyond this point are not substantial enough to justify the higher resource demands.
Study on the mask selection strategy. To investigate the impact of mask selection strategies on final segmentation performance, we experimented with different strategies. As shown in Table 7, we first attempted to select a single mask as the final segmentation result by choosing the mask with the highest $S_{sim}$ or $S_{iop}$ score. Experimental results show that this single-mask selection strategy performs significantly worse than the multi-mask selection strategy. Therefore, among the five strategies, we choose one that considers masks with high semantic similarity to the text features while also selecting as many masks as possible with high $S_{iop}$ scores. This strategy achieved gIoU of 56.9 and cIoU of 54.4 on the ReasonSeg validation set. By adopting this selection strategy, HRSeg effectively combines multiple over-segmented proposals from SAM, leading to optimal performance. A qualitative analysis of this strategy can be found in the supplementary materials. It demonstrates how our strategy works in complex scenarios, and compared to other models, the final segmentation performance is notably improved due to the high-resolution perception technique introduced by HRSeg.
Furthermore, we conducted experiments on the threshold of $S_{iop}$. As shown in Table 8, the initial threshold of 0.5 did not yield the best performance. Instead, the optimal performance was achieved when the threshold was set to 0.7. This suggests that the initial setting included some low-confidence masks, which negatively impacted segmentation performance.
@@ 4.7. Qualitative Results
Fig. 4 presents visualization examples from the ReasonSeg dataset, comparing HRSeg with other SOTA methods. It is evident that HRSeg consistently delivers more accurate and detailed segmentation results. As shown in the first three rows of the figure, HRSeg performs exceptionally well in scenarios involving small objects or situations that require more detailed visual perception. We attribute this to the key components of HRSeg, which enhance its ability to perceive fine-grained visual details, enabling it to interpret implicit instructions and achieve precise segmentation. In contrast, other methods often produce incorrect or imprecise segmentation results in complex scenarios. More visualization results, including analyses of how the model perceives fine-grained details and its failure cases, can be found in the supplementary materials.

### v2

@@ Abstract.
The reasoning segmentation task involves segmenting objects within an image by interpreting implicit user instructions. Despite significant advancements made by existing approaches, they remain constrained by low perceptual resolution, as visual encoders are typically pre-trained at lower resolutions. Furthermore, simply interpolating the positional embeddings of visual encoders to enhance perceptual resolution yields only marginal performance improvements while incurring substantial computational costs. To address this, we propose HRSeg, an efficient model with high-resolution fine-grained perception. It features two key innovations: High-Resolution Perception (HRP) and High-Resolution Enhancement (HRE). The HRP module processes high-resolution images through cropping, integrating local and global features for multi-granularity quality. The HRE module enhances mask features by integrating fine-grained information from high-resolution images, refining their alignment with text features for precise segmentation. Extensive ablation studies validate the effectiveness of our modules, while comprehensive experiments on multiple benchmark datasets demonstrate HRSeg’s superior performance. Code will be available at https://github.com/WeihuangLin/HRSeg.
Reasoning Segmentation; Multimodal Large Language Models; High-resolution Perception
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755006isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Image segmentation
@@ 3.1. Overview
Method val test
overall short query long query overall
gIoU cIoU gIoU cIoU gIoU cIoU gIoU cIoU
OVSeg (Liang et al., 2023) 28.5 18.6 18.0 15.5 28.7 22.5 26.1 20.8
GRES (Liu et al., 2023a) 22.4 19.9 17.6 15.0 22.6 23.8 21.3 22.0
X-Decoder (Zou et al., 2023) 22.6 17.9 20.4 11.6 22.2 17.5 21.7 16.3
SEEM (Zou et al., 2024) 25.5 21.2 20.1 11.5 25.6 20.8 24.3 18.7
LISA (Lai et al., 2024) 44.4 46.0 37.6 34.4 36.6 34.7 36.8 34.1
LLM-Seg (Wang and Ke, 2024) 47.4 35.4 39.8 32.9 41.3 33.8 40.2 33.4
SAM4MLLM (Chen et al., 2024a) 46.7 48.1 - - - - - -
HRSeg 50.1 47.1 42.1 37.7 47.9 43.7 46.6 42.6
HRSeg-LLaVA-v1.6 56.7 55.4 47.9 44.2 56.6 53.2 54.6 51.2
LISA (ft) 51.0 50.6 40.6 40.6 49.4 51.0 47.3 48.4
LLM-Seg (ft) 52.3 47.5 41.1 40.2 49.8 49.1 47.9 46.2
HRSeg (ft) 57.4 54.7 43.4 41.7 54.4 54.5 51.6 50.5
HRSeg-LLaVA-v1.6 (ft) 64.9 63.1 49.5 48.7 59.2 58.9 57.0 57.2
@@ 4.2. Implementation Details
We trained HRSeg for one day using the DeepSpeed engine. The mask proposals are generated using a $32\times 32$ grid of points to prompt SAM. The default MLLM is LLaVA-7B-v1-1, with LoRA applied for efficient fine-tuning. The image encoder used for the HRP module is DINOv2-ViT-L (Oquab et al., 2023), which is pretrained on images of size $518\times 518$ pixels. We use AdamW (Loshchilov, 2017) as the optimizer and apply WarmupDecayLR to adjust the learning rate. The initial learning rate is set to 0.0003. We use a batch size of 2 per GPU. For high-resolution images, we set $N=2$ by default, which corresponds to an image size of $518\times 2=1036$ . The $\gamma$ in the HRP module is set to 0.8 to weight the features, and the $L$ value is set to 2 in the HRE module. For the loss function, the weights for the text loss ($\lambda_{text}$) and selection loss ($\lambda_{sel}$) are both set to 1.0.
@@ 4.3. Results on ReasonSeg
Table 1 provides a detailed comparison between our proposed HRSeg and several state-of-the-art segmentation models. HRSeg is specifically designed for high-resolution reasoning segmentation tasks.
Experimental results show that HRSeg consistently outperforms existing methods across all metrics on the ReasonSeg dataset. Without fine-tuning, it surpasses LLM-Seg by 2.7%/8.8% (gIoU/cIoU) on the validation set and 6.4%/9.2% on the test set. HRSeg also handles both short and long queries better, with respective gains of 5.6%/11.0% and 6.6%/9.9%. Fine-tuning on the ReasonSeg training split further enhances HRSeg’s performance across all evaluation metrics.
In particular, for long-query instructions that require richer visual understanding, HRSeg outperforms LLM-Seg by 6.6%, which is 4.3% higher than the improvement in short-query scenarios. This indicates HRSeg’s ability to capture fine-grained visual features, enabling high-quality mask generation and better alignment with complex queries.
We attribute HRSeg’s superior performance to its use of high-resolution visual information, which improves both segmentation detail and the model’s ability to handle implicit textual cues, ultimately boosting overall accuracy.
Method refCOCO refCOCO+ refCOCOg
val testA testB val testA testB val(U) test(U)
MCN (Luo et al., 2020) 62.4 64.2 59.7 50.6 55.0 44.7 49.2 49.4
VLT (Ding et al., 2021) 67.5 70.5 65.2 56.3 61.0 50.1 55.0 57.7
CRIS (Wang et al., 2022) 70.5 73.2 66.1 62.3 68.1 53.7 59.9 60.4
LAVT (Yang et al., 2022) 72.7 75.8 68.8 62.1 68.4 55.1 61.2 62.1
ReLA (Liu et al., 2023a) 73.8 76.5 70.2 66.0 71.0 57.7 65.0 66.0
HRSeg-LLaVA-v1 77.1 79.7 73.2 66.4 72.4 58.9 71.4 72.8
HRSeg-LLaVA-v1.6 81.2 83.4 78.1 73.7 79.1 67.1 75.6 76.4
@@ 4.6. Ablation Study
Study on the weight $\gamma$. To assess the effect of weighting global and region features, we analyze different values of $\gamma$ in the HRP module. As shown in Table 5, $\gamma$ significantly impacts the model’s ability to integrate these feature types. A low $\gamma$ weakens global context modeling, while a high value suppresses crucial region-specific details. The optimal setting, $\gamma=0.8$, strikes a balance, effectively combining global and local cues to improve reasoning segmentation performance.
@@
Study on the image resolution. To further investigate the impact of increased image resolution on the performance of our model, we conducted an ablation study on image resolution. As shown in Table 6, $N$ denotes the magnification factor relative to the pre-trained encoder’s input. Results show that higher resolutions improve segmentation by capturing finer details, benefiting both global and local reasoning. However, the performance gains taper off as resolution increases, while computational cost rises sharply. Therefore, we select $N=2$ as the optimal setting, offering a favorable trade-off between accuracy and efficiency.
Study on the mask selection strategy. To investigate the impact of mask selection strategies on final segmentation performance, we experimented with different strategies. As shown in Table 7, we first evaluated single-mask selection using either the highest $S_{sim}$ or $S_{iop}$ score, but this approach performed significantly worse than multi-mask strategies. Among the five options, the most effective strategy considers masks with high semantic similarity to text features while maximizing those with high $S_{iop}$ scores. This approach achieved 56.9 gIoU and 54.4 cIoU on the ReasonSeg validation set. By combining multiple over-segmented SAM proposals, this strategy enables HRSeg to reach optimal performance. Furthermore, we conducted experiments on the threshold of $S_{iop}$. As shown in Table 8, the initial threshold of 0.5 did not yield the best performance. Instead, the optimal performance was achieved when the threshold was set to 0.7. This suggests that the initial setting included some low-confidence masks, which negatively impacted segmentation performance.
@@ 4.7. Qualitative Results
Fig. 4 shows qualitative results on the ReasonSeg dataset, comparing HRSeg with other SOTA methods. HRSeg consistently produces more accurate and detailed segmentations. In the first three rows, it performs especially well on small objects and visually demanding scenarios. This is attributed to HRSeg’s key components that enhance fine-grained perception, allowing it to handle implicit instructions and achieve precise segmentation. In contrast, other methods often fail in complex cases, yielding inaccurate or coarse results.
@@ 5. Conclusion
Acknowledgements.
This work was supported by National Key R&D Program of China (No.2023YFB4502804), the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U22B2051, No. U21B2037, No. 62302411, No. 624B2118), the Natural Science Foundation of Fujian Province of China (No.2021J06003), and China Postdoctoral Science Foundation (No. 2023M732948).